The speed of an ordinary train is x km/hr and that of an express train is x+25 km/hr. (i) Find the time taken for each train to cover 300 km. (ii) If the ordinary tran takes 2 hrs more than the express train; calculate the speed of the express train.
Answer by TimothyLamb(4379)

s = d/t
t = d/s
---
time for ordinary train:
a = 300/x
---
time for express train:
b = 300/(x+25)
---
a = b + 2
300/x = 300/(x+25) + 2
300/x - 300/(x+25) = 2
(x+25)300/x(x+25) - 300x/x(x+25) = 2
(x+25)300 - 300x = 2x(x+25)
300x + 25*300 - 300x = 2xx + 50x
2xx + 50x - 7500 = 0
---
the above quadratic equation is in standard form, with a=2, b=50 and c=-7500
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 50 -7500
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 50
x = -75
---
the negative root doesn't fit the problem statement, so use the positive root:
x = 50
---
answer:
speed of the express train = 50 + 25 = 75 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php